OpenStudy (anonymous):

List steps - subtract - simplify -5/10 - 5/10

OpenStudy (anonymous):

-5-5=-10 isn't it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-10/10 what it is?

OpenStudy (anonymous):

-5 -5 = -10 -10/10 = -1

OpenStudy (lalaly):

\[\frac{-5-5}{10}= \frac{-10}{10}=-1\]

OpenStudy (anonymous):

oh its -1! i didn't know that :/

OpenStudy (anonymous):

thank yous very much